OpenStudy (anonymous):

intergral (r^3)/sqrt(4+r^2) from 0 to 1

OpenStudy (anonymous):

\[\int_0^1\frac{r^3}{\sqrt{4+r^2}}dr\] Have you tried the following substitution? \[r=2\tan u\]

OpenStudy (anonymous):

yes i did

OpenStudy (anonymous):

\[dr=2\sec^2 u \;du\\ \text{The integral then becomes}\\ \int_0^{2\tan1}\dfrac{(2\tan u)^2(2\sec^2u)}{\sqrt{4+(2\tan u)^2}}du\] What can you do with that? And if you have any questions about how you get the above, ask away.

OpenStudy (anonymous):

Oh, I've made a slight mistake. In the numerator you should have \[(2\tan u)^3 \text{ instead of }(2 \tan u)^2.\]

OpenStudy (anonymous):

how did you get the bounds of the integral

OpenStudy (anonymous):

Ah, that's the second time I've made that mistake. They should be . Sorry! Since r = 2 tanu, we know that u = arctan(r/2) Upper u = arctan(1/2) Lower u = arctan(0/2) = arctan(0) = 0 So the integral should be fixed now, \[\int_0^{\arctan2}\frac{(2\tan u)^3 (2\sec^2u)}{\sqrt{4+(2\tan u)^2}}du\]

OpenStudy (anonymous):

**They should be 0 and arctan(1/2)**

OpenStudy (anonymous):

i get 2^3(1/3(sec^3u)-secu)

OpenStudy (anonymous):

as my derivative of the intergral

OpenStudy (anonymous):

What do you mean by "derivative of the integral"? You should be getting a number as your answer.

OpenStudy (anonymous):

before i minus the lower bound from the upper bound

OpenStudy (anonymous):

my final answer i get (1/3)((-7sqrt5) +16)

OpenStudy (anonymous):

I don't see where the 1/3 is coming from. Here's what you should be getting, step by step. \[\int_{0}^{\arctan2}\frac{2^4 \tan^3u\sec^2u}{\sqrt{4+4\tan^2u}}du\\ 16\int_{0}^{\arctan2}\frac{\tan^3u\sec^2u}{\sqrt{4}\sqrt{1+\tan^2u}}du\\ 8\int_{0}^{\arctan2}\frac{\tan^3u\sec^2u}{\sqrt{\sec^2u}}du\\ 8\int_{0}^{\arctan2}\frac{\tan^3u\sec^2u}{\sec u}du\\ 8\int_{0}^{\arctan2}\tan^3u\sec udu\\ 8\int_{0}^{\arctan2}\tan^2u\sec u\tan udu\\ 8\int_{0}^{\arctan2}(\sec^2u-1)\sec u\tan udu\\ 8\int_{0}^{\arctan2}\left[\sec^2u\sec u\tan u-\sec u\tan u\right]du\\ 8\left[\int_{0}^{\arctan2}\sec^2u\sec u\tan u\;du-\int_{0}^{\arctan2}\sec u\tan u\;du\right]\] For the first integral, let t = secu, dt = secu tanu du \[8\left[\int_{1}^{\sqrt5}t^2\;dt-\int_{0}^{\arctan2}\sec u\tan u\;du\right]\] The rest is pretty easy.

OpenStudy (anonymous):

i used 8*intergral of t^2-1 dt

OpenStudy (anonymous):

when i intergrate that i get 1/3t^3-t

OpenStudy (anonymous):

Okay. What about the 1, in (t² - 1)? Is that what secu tanu becomes?

OpenStudy (anonymous):

doesnt 1 become t when you int. and t=secu

OpenStudy (anonymous):

Not quite. The second integral, \[\int_0^{\arctan2}\sec u\tan u \;du = \left[\sec u\right]_0^{\arctan2}\] If you were to use the substitution t = secu, you'd have \[[t]_1^{\sqrt5}\] Besides, it's not clear how you get t²-1 in the first place. You have the int(t² dt) - int(secu tanu du), and neither of these terms reduce to 1.